OpenStudy (anonymous):

If y+3=11,what is the value of y–7?

OpenStudy (anonymous):

y equals 8, so 8 -7 =1

OpenStudy (anonymous):

11-7=y

OpenStudy (anonymous):

for the first eqation 11 - 3 =8 you use 8 and plug it in the second equation

OpenStudy (anonymous):

11-7=y y-7=1 7+1=8=y

OpenStudy (solomonzelman):

I see you are still confused because the question is not closed, and you are still viewing it. I am not good at explaining, but I'll take a shot anyway. If y+3=11,what is the value of y–7? If y+3=11 lets solve for y, y+3=11 -3 -3 y = 8 Okay? so we got y=8 now, what is y-7? saying y-7 is the same as saying 8-7, knowing that y=8. hence is your answer, (8-7 equals) 1.

OpenStudy (anonymous):

Y = 8 because you would subtract 3 from 11 which would be 8.